Uber has a monopoly on ride-sharing services. In one town, the demand curve on weekdays is given by the following equation: P = 50 - Q. However, during weekend nights, or peak hours, the demand for rides increases dramatically and the new demand curve is P = 100 - Q. Assume that marginal cost is zero.


a. Determine the profit-maximizing price during weekdays and during peak hours.

b. Determine the profit-maximizing price during weekdays and during peak hours if MC = 10 instead of zero.

c. Draw a graph showing the demand, marginal revenue, and marginal cost curves during peak hours from part (b), indicating the profit-maximizing price and quantity. Determine Uber’s profit and the deadweight loss during peak hours, and show them on the graph.


1
Expert's answer
2022-03-23T14:02:28-0400


Total revenue during weekdays

"Price \times quantity"

"PQ = (50-Q)Q = 50Q- Q^2"

Marginal revenue is

"\frac{\Delta TR}{\Delta Q}=50-2Q"

"Q=25"

"P = 25"


During peak hours

"PQ= (100-Q)Q"

"= 100Q-Q^2"

"MR=100-2Q"

"MC =0"

"Q= 50"

"P= 50"

When Marginal cost changes to 10

Weekdays;

"50-2Q =10"

"Q=20"

"P=30"

Weekends and peak hours

"100-2Q =10"

"Q= 45"

"P= 55"